Hi, I am having a REALLY hard time with word problems right now. Could someone PLEASE help me with this problem? Help would be greatly appreciated, thanks! Here is the problem: A Garden plot 4 m by 12 m has one side along a fence as shown at the right. The area of the garden is to be doubled by digging a border of uniform width on the other three sides. What should the width of the border be?
Please HELP me??? Thank you so much!

Answer by stanbon(75887)

Draw the picture of the original garden with
width = 4 and length = 12
Draw a uniform border around three sides having
a uniform width of "x".
The new width is now 4+x; the new length is now 12 + 2x
The new area is (4+x)(12+2x) = 96
48+20x+2x^2 = 96
2x^2+20x-48=0
x^2+10x-24=0
(x+12)(x-2)=0
Therefore x=2 meters is the uniform width of the border.
Cheers,
Stan H.